OpenStudy (anonymous):

Heisenburgs Uncertainty Principle. I keep getting the wrong answer and it is driving me crazy. Using Heisenbergs uncertanty priciple, calculate the uncertanty in the position of a) A 1.50-mg mosquito moving at a speed of 1.40 m/s if the speed is known to within 0.01 m/s; and b) which I will try to figure out on my own if I fully understand this one. The following post will show what I have done.

OpenStudy (anonymous):

\[h= 6.626x10^{-34}\] \[\Delta x = h/(4\pi *\Delta p) \] \[m = 1.5mg = 1.5 * 10^{6}\] \[\Delta v = 1.4 m/s \pm0.01 m/s = 1.40*10^{-2}\]

OpenStudy (anonymous):

I would assume, here that I have everything and it should just be plug and chug, however I am not getting the correct answer and do not know why.

OpenStudy (jfraser):

1.5mg isn't 1.5x10^6 kg, it's 1.5x10^-6 kg. You're off by 12 orders of magnitude

OpenStudy (anonymous):

That was a typo, the magnitude is not what I am worried about, it is the actual value.

OpenStudy (anonymous):

Answered in http://openstudy.com/study#/updates/509ebd14e4b013fc35a17ac3